# SincVAE: a New Approach to Improve Anomaly Detection on EEG Data Using SincNet and Variational Autoencoder

### v1

@@ 6 Conclusions
Also, the capability of SincVAE to discern subtle anomalies in EEG data indicates SincVAE as a tool to detect early signs of epilepsy in the preictal stage and to monitor the patients status during the postictal stage. This aspect can profoundly affect patient monitoring.
Furthermore, the semi-supervised nature of SincVAE, requiring only non-seizure data for training, addresses challenges associated with the scarcity of labeled anomaly data in medical datasets, making it an efficient solution in real-world applications where anomalies are rare but critical to detect.
Future work could explore the application of SincVAE in other types of time-series anomaly detection tasks, potentially broadening its utility in healthcare and other fields.
@@ Acknowledgement
This work was supported by PRIN research project PE0000013 Future Artificial Intelligence Research – FAIR, E63C22002150007

### v2

@@ 6 Conclusions
Also, the capability of SincVAE to discern subtle anomalies in EEG data indicates SincVAE as a tool to detect early signs of epilepsy in the preictal stage and to monitor the patients’ status during the postictal stage. This aspect can profoundly affect patient monitoring.
Furthermore, the semi-supervised nature of SincVAE, requiring only non-seizure data for training, addresses challenges associated with the scarcity of labeled anomaly data in medical datasets, making it an efficient solution in real-world applications where anomalies are rare but critical to detect. However, it is important to note that, like any DL-based method, SincVAE is data-driven and its performance heavily depends on the availability of extensive, high-quality data. A comprehensive dataset that adequately represents the patient’s normal state is essential for ensuring the model generalizes effectively, enabling it to reliably differentiate between normal and anomalous conditions.
Future work could explore the application of SincVAE in other types of time-series anomaly detection tasks, potentially broadening its utility in healthcare. Another promising direction involves integrating attention mechanisms [117] within the SincVAE framework, in particular on the SincNet layer, to refine the selection of frequency bands identified by the SincNet layer. Such enhancement could enable more accurate filtering by prioritizing certain frequency bands more than others, potentially leading to an improvement of the model’s overall performance.
Moreover, a key direction for future work will be enhancing the interpretability of the SincNet layer to ensure greater transparency and provide deeper insights, particularly within clinical contexts. As illustrated in Figures 15 and 16, the learned filters and corresponding filtered signals become observable and interpretable after training. Attention coefficients may offer valuable insights into the relative importance of different filters, thereby providing a clearer understanding of which frequency bands are most influential for the task at hand. Also, we intend to explore the use of explainable artificial intelligence (XAI) [118, 119] techniques in this direction. These methods could provide additional interpretative power, allowing for a deeper analysis of how the model processes and prioritizes different frequency bands. Finally, we also plan to conduct ablation studies [120, 121] to systematically evaluate the contribution of specific frequency bands to the model’s performance. These studies will be crucial in identifying the most reliable direction to improve interpretability of the filtering stage of the SincNet layer, in order to let the analysts understand which frequency bands have the most significant impact on the solution with respect to the input data, thereby assisting in interpreting the outcomes from a clinical perspective.
@@ Acknowledgement
This work was supported by the Italian Ministry of University and Research, PNRR project PE00000013 Future Artificial Intelligence Research (FAIR), E63C22002150007 and by the Italian Ministry of University and Research, PRIN research project “BRIO – BIAS, RISK, OPACITY in AI: design, verification and development of Trustworthy AI.”, Project no. 2020SSKZ7R. The authors express their gratitude to Dr. Angela Natalizio and Dr. Antonio Esposito for their support in enhancing the understanding of the biomedical aspects of the analyzed signals.